OpenStudy (anonymous):

Simplify The Expression 4^3 ⋅ 4^5

OpenStudy (anonymous):

65536

OpenStudy (anonymous):

4^3 is the same thing as 4*4*4 4^5 is the same thing as 4*4*4*4*4 So 4^3 * 4^5 is the same as 4*4*4*4*4*4*4*4 Which is equal to 4^8 So whenever you multiply with the same base(4 here) you can simply add up the exponents.

OpenStudy (solomonzelman):

\[4^3 \times 4^5=4^{3+5}=4^8\]

OpenStudy (solomonzelman):

See how I am adding exponents when multiplying, @freakenpurplepizza ?